OpenStudy (anonymous):

x-y=-2 -x-8y=2 Solve With Each System Of Elimination

imqwerty (imqwerty):

Try adding the two equations :) what do u get??

OpenStudy (anonymous):

imqwerty Please Help Again I Love How U Do It

OpenStudy (anonymous):

-2,0 Is The Answer I Got

imqwerty (imqwerty):

Correct!! :D

OpenStudy (anonymous):

Im Gonna Fan You .. But Can You Show Me Steps

imqwerty (imqwerty):

Ok So we have - x-y=-2 1st equation -x-8y=2 2nd equation So when we add both these equations we get- x-y +(-x-8y)=-2+2 x-y-x-8y=0 -9y=0 y=0/-9 y=0 Puttin y=0 in the first we get x-y=-2 x-0=-2 x=-2

OpenStudy (anonymous):

You Helped Alot !